OpenStudy (anonymous):

A taxi service charges an initial fee of $1.50 plus $1.50 for every mile traveled. A taxi ride costs $21.00. How many miles did the taxi travel?

OpenStudy (paki):

1.50+1.50 = 3$ per mile... now here for 21$, it will be 21/3 = ??? miles

OpenStudy (mathstudent55):

cost = initial fixed fee + per mile cost The cost per mile is $1.50. That means: 1 miles costs 1.50 * 1 2 miles cost 1.50 * 2 3 miles cost 1.50 * 3 x miles cost 1.50 * x = 1.5x cost = initial fixed fee + 1.5x Now substitute 21 for cost, and 1.50 for the initial fixed fee. Then solve the equation for x.

OpenStudy (mathstudent55):

@paki Be careful. The way you explained it, you are charging the initial fee for every mile. The initial fee is only charged once. The number of miles is definitely more than 7.

OpenStudy (anonymous):

thank you

OpenStudy (mathstudent55):

You're welcome.